OpenStudy (anonymous):

Draw a line with a slope of 1/4 through the point (0, -4).

OpenStudy (anonymous):

A:

OpenStudy (anonymous):

B:

OpenStudy (anonymous):

C:

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

Do you know what slope-intercept form is?

OpenStudy (anonymous):

I forgot My brain is not working today

OpenStudy (anonymous):

I really need the answer

OpenStudy (anonymous):

y = mx + b, where m is the slope, which we have

OpenStudy (anonymous):

So what again is the slope?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok now its C

OpenStudy (anonymous):

The slope is 1/4, so the equation is now y = 1/4x + b

OpenStudy (anonymous):

b is where the graph intercepts the y axis. We can tell when a coordinate is the y intercept when the x value is also 0

OpenStudy (anonymous):

So since we have (0, -4), that means the y intercept is?

OpenStudy (anonymous):

Nvm, i already got the answer nut you still get a medal tho

OpenStudy (anonymous):

*but not nut

OpenStudy (anonymous):

lol